Assuming 100% dissociation, calculate the freezing point and boiling point of 2.97 m AgNO3(aq).
Answer #1

dTf = -Kf *m * i

i = 2 ions (ag+ and NO3-)

m = 2.97

Kf = 1.86

then

dTf = -1.86*2.97*2 = -11.048

TF = -11.048 °C